OpenStudy (anonymous):

a rectangular box has a square base. The volume of the box is 256 cubic centimeters. The material for the base and the top of the box cost 0.10 per square centimenter and the material for the sides cost 0.05 per square centimenter. Find the dimensions of the box that minimize the cost of producing box.

OpenStudy (anonymous):

no wonder you really hate math

OpenStudy (anonymous):

LOOL calculus kills me

OpenStudy (anonymous):

we can do this if you want

OpenStudy (anonymous):

omg you know how to do that :)

OpenStudy (anonymous):

yeah it will just take a few minutes. ok base is a square label the side x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so area of the bottom and top are both \[x^2\]

OpenStudy (anonymous):

they cost .10 per square whatever and you will have two of them (top and bottom) so total cost for the top and bottom is \[.10\times 2x^2\]

OpenStudy (anonymous):

or if you prefer \[.2x^2\] these decimals are killing me. ok we continue

OpenStudy (anonymous):

now the volume is 256 which we get by multiplying area of base times height.

OpenStudy (anonymous):

the base has area \[x^2\] so we know \[x^2h=256\] i.e. \[h=\frac{256}{h}\]

OpenStudy (anonymous):

sorry \[h=\frac{256}{x}\]

OpenStudy (anonymous):

still with me?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

ok we have 4 sides. the base of each side is x and the height is \[h=\frac{256}{x^2}\] so their area is each \[x\times \frac{256}{x^2}=\frac{256}{x}\]

OpenStudy (anonymous):

there are four of them so multiply by 4 to get the total area of the sides all together

OpenStudy (anonymous):

i get 1024/x

OpenStudy (anonymous):

so we multiply that by .05

OpenStudy (anonymous):

and these each cost .05 (dollars, whatever)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get 51.2/x

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

now i forget what the other stuff cost.

OpenStudy (anonymous):

oh .2x^2

OpenStudy (anonymous):

ok so total cost is\[ .2x^2+\frac{51.2}{x}\]

OpenStudy (anonymous):

and this is the damned thing you have to find the minimum of. now for the easy part. take the derivative, find the critical points etc. ready?

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

actually it is not that bad.

OpenStudy (anonymous):

\[\frac{d}{dx} .2x^2+\frac{51.2}{x}=.4x-\frac{51.2}{x^2}\]

OpenStudy (anonymous):

now set = 0 and solve \[.4x-\frac{51.2}{x^2}=0\] \[\frac{.4x^3-51.2}{x^3}=0\] \[.4x^3-51.2=0\]

OpenStudy (anonymous):

\[x^3=\frac{51.2}{.4}=128\]

OpenStudy (anonymous):

and finally..... \[x=\sqrt[3]{128}\]

OpenStudy (anonymous):

5.04 rounded. i am exhausted!

OpenStudy (anonymous):

omg I think I love u haha

OpenStudy (anonymous):

please please dont tell me the answer in the back of the book is 12!

OpenStudy (anonymous):

thanx radar!

OpenStudy (anonymous):

looooll no

OpenStudy (anonymous):

is that all or do you have a pile of these wonders to work out?

OpenStudy (anonymous):

I have one more lool but I dont wanna bother you again

OpenStudy (anonymous):

go ahead if i get tired i will leave

OpenStudy (anonymous):

sqroot (x+y)=1+((x^2)(y^2))

OpenStudy (anonymous):

ahh implicit diff?

OpenStudy (anonymous):

your teacher must hate you

OpenStudy (anonymous):

find the y' and compute the value at (0,1)

OpenStudy (anonymous):

ok we differentiate wrt x

OpenStudy (anonymous):

remembering that when we take the derivative of y we get y'

OpenStudy (anonymous):

here goes \[\sqrt{x+y}=1+x^2y^2\]

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{x+y}}\times y'=2xy^2+x^22yy'\]

OpenStudy (anonymous):

left hand side was because the derivative of root x is one over two root x, and the y' by the chain rule. left hand side was the product rule. first the derivative of x^2 is 2x and next the derivative of y^2 is 2y y' again by the chain rule

OpenStudy (anonymous):

now we solve for y' \[\frac{y'}{\sqrt{x+y}}-2x^2yy'=2xy^2\] \[y'(\frac{1}{\sqrt{x+y}}-2x^2y)=2xy^2\] \[y'=\frac{2xy^2}{\frac{1}{\sqrt{x+y}}-2x^2y}\]

OpenStudy (anonymous):

i forgot the 2 outside the square root sign. sorry. it is there.

OpenStudy (anonymous):

i will simplify that if you like

OpenStudy (anonymous):

where are the sqroot lol

OpenStudy (anonymous):

in the bottom

OpenStudy (anonymous):

here it is simplified

OpenStudy (anonymous):

\[\frac{4xy^2\sqrt{x+y}-1}{1-4x^2y\sqrt{x+y}}\]

OpenStudy (anonymous):

now plugging in (0,1) should be nice and easy so i will let you do that. have fun!

OpenStudy (anonymous):

loool thx a LOT good nite:)